# A Statistical Viewpoint on Differential Privacy: Hypothesis Testing, Representation and Blackwell’s Theorem

### v1

@@ Abstract
Differential privacy is widely considered the formal privacy for privacy-preserving data analysis due to its robust and rigorous guarantees, with increasingly broad adoption in public services, academia, and industry. Despite originating in the cryptographic context, in this review paper we argue that, fundamentally, differential privacy can be considered a pure statistical concept. By leveraging a theorem due to David Blackwell, our focus is to demonstrate that the definition of differential privacy can be formally motivated from a hypothesis testing perspective, thereby showing that hypothesis testing is not merely convenient but also the right language for reasoning about differential privacy. This insight leads to the definition of $f$-differential privacy, which extends other differential privacy definitions through a representation theorem. We review techniques that render $f$-differential privacy a unified framework for analyzing privacy bounds in data analysis and machine learning. Applications of this differential privacy definition to private deep learning, private convex optimization, shuffled mechanisms, and U.S. Census data are discussed to highlight the benefits of analyzing privacy bounds under this framework compared to existing alternatives.
@@ 1 Introduction
[DRS22] formally addressed this question in the affirmative. The core of their strategy is to leverage a result due to David Blackwell to identify a measure that encodes the complete information for hypothesis testing in the context of differential privacy [Bla50]. Roughly speaking, this information-complete measure is given by the envelope of all uniformly most powerful tests using the likelihood ratio, as characterized by the Neyman–Pearson lemma. With this measure in place, [DRS22] introduced $f$-differential privacy and showed, using a representation theorem, that any differential privacy definition could be considered a subfamily of this privacy definition.
@@ 2.2 Hypothesis Testing Interpretation
Having introduced the background that motivated differential privacy, in this section we formally define this privacy notion. The “differential” nature of differential privacy can be precisely formulated as a hypothesis testing problem in which an attacker makes a choice between [WZ10]
@@
For this hypothesis testing problem, we envision a powerful attacker who knows both the output of the algorithm and its probability distributions under both the null $H_{0}$ and alternative $H_{1}$. This probability distribution fully characterizes the information provided by the algorithm. By assuming such a strong attacker, we ensure that the resulting privacy definition, which can render the attacker powerless, is inherently strong.
Having formulated differential privacy from the hypothesis testing perspective, the next step is to specify the basis upon which the attacker determines which hypothesis to reject or accept. This necessitates a metric between the two probability distributions corresponding to the null and the alternative to quantify their difference. This motivates us to propose the following axiom as a self-evident requirement for any possible differential privacy definitions.
@@
Formally, denote by $\mathrm{Proc}$ any post-processing operation on the output of the algorithm, such as dimension reduction or summarization. Axiom 2 states that $D(M(S),M(S^{\prime}))$ and $D(\mathrm{Proc}(M(S)),\mathrm{Proc}(M(S^{\prime})))$ must be comparable and satisfy
@@ 2.3 A Representation Theorem
A proof of Theorem 3 is given in Section 2.4. For the construction of link functions for certain types of metrics $D$, see Proposition B.4 in [DRS22].
@@ 1. Pure differential privacy
    to define pure differential privacy. The corresponding link function is $d_{\infty}(f)=-\log(-f^{\prime}(1))$, where the derivative is understood as the left derivative. It is possible that $f^{\prime}(1)=0-$. Thus, $\epsilon=\infty$ is included for completeness.
@@ 2. Approximate differential privacy
    where $x_{+}$ denotes $\max\{x,0\}$. An algorithm $M$ satisfies $(\epsilon,\delta)$-differential privacy if $D_{\infty}^{\delta}(M(S),M(S^{\prime}))$ is upper bounded by $\epsilon$ for any neighboring datasets $S$ and $S^{\prime}$. Smaller values of $\epsilon$ and $\delta$ provide stronger privacy protection. More precisely, $\delta$ should be set to a very small value, while a relatively large value of $\epsilon$ can still provide meaningful privacy guarantees [DKM+06, ND22].
@@
    where $f^{\ast}(x)=\sup_{y}(xy-f(y))$ is the convex conjugate of $f=T(P,Q)$. By varying $\delta$, all $(\epsilon,\delta)$-differential privacy guarantees taken together give the privacy profiles of the algorithm [BBG18], which are the dual representation of the trade-off function $f$.
@@ 3.2 Achieving Differential Privacy
[AV23] called $Z$ a canonical noise distribution for $f$ if the three properties above are satisfied. Notably, a canonical noise distribution fully exploits the privacy budget conveyed by the trade-off function. This distribution admits an explicit construction based on the given trade-off function [AV23]. It has been extended to multivariate settings and shown to attain certain optimal statistical properties [AD22, AR23].
@@ 4.3 Shuffled Mechanism
The DAS utilizes a sequence of counting queries $\left\{\widebar{M}_{i}(S)\right\}_{i=1}^{m}$ with $m=72$, as specified in the privacy allocation file [U.S22]. These queries cover 8 geographical levels, including nation, state, and county, with each level containing 9 queries. Each query is privatized as $M_{i}(S)=\widebar{M}_{i}(S)+\mathcal{N}_{\mathbb{Z}}(0,\sigma_{i}^{2})$. As each $\widebar{M}(S)$ is a counting query with sensitivity 1, the analysis of the privacy bound boils down to evaluating the trade-off function for testing the following hypotheses:
@@ 5 Concluding Remarks
In this paper, we have explored differential privacy from a statistical viewpoint, motivated by the hypothesis testing framework. We have shown that hypothesis testing is not merely a convenient tool but in fact the right language for reasoning about differential privacy. In light of this, $f$-differential privacy becomes a very natural definition that preserves both the hypothesis testing interpretation and the property of post-processing. Owing to its coherent mathematical structure, the $f$-differential privacy framework is accompanied by powerful tools, such as composition theorems, which enable tight and even lossless privacy analysis. This paper has demonstrated the effectiveness of this privacy definition in several important applications. Nevertheless, the review does not cover other notable applications, including federated learning, privacy auditing, and Markov chain Monte Carlo [ZCLS21, NHS+23, LMWR23].

### v2

@@ Abstract
Differential privacy is widely considered the formal privacy for privacy-preserving data analysis due to its robust and rigorous guarantees, with increasingly broad adoption in public services, academia, and industry. Despite originating in the cryptographic context, in this review paper we argue that, fundamentally, differential privacy can be considered a pure statistical concept. By leveraging David Blackwell’s informativeness theorem, our focus is to demonstrate based on prior work that all definitions of differential privacy can be formally motivated from a hypothesis testing perspective, thereby showing that hypothesis testing is not merely convenient but also the right language for reasoning about differential privacy. This insight leads to the definition of $f$-differential privacy, which extends other differential privacy definitions through a representation theorem. We review techniques that render $f$-differential privacy a unified framework for analyzing privacy bounds in data analysis and machine learning. Applications of this differential privacy definition to private deep learning, private convex optimization, shuffled mechanisms, and U.S. Census data are discussed to highlight the benefits of analyzing privacy bounds under this framework compared to existing alternatives.
@@ 1 Introduction
In [KOV17, DRS22], the authors formally addressed this question in the affirmative. The core of their strategies is to leverage a result due to David Blackwell to identify a measure that encodes the complete information for hypothesis testing in the context of differential privacy [Bla50]. Roughly speaking, this information-complete measure is given by the envelope of all uniformly most powerful tests using the likelihood ratio, as characterized by the Neyman–Pearson lemma. With this measure in place, [DRS22] introduced $f$-differential privacy and showed, using a representation theorem, that any differential privacy definition could be considered a subfamily of this privacy definition.
@@ 2.2 Hypothesis Testing Interpretation
Having introduced the background that motivated differential privacy, in this section we formally define this privacy notion. The “differential” nature of differential privacy can be precisely formulated as a hypothesis testing problem in which an attacker makes a choice between
@@
This connection was first recognized by [WZ10] and further refined by [KOV17]. To interpret this hypothesis testing problem, we envision a powerful attacker who knows both the output of the algorithm and its probability distributions under both the null $H_{0}$ and alternative $H_{1}$. This probability distribution fully characterizes the information provided by the algorithm. By assuming such a strong attacker, we ensure that the resulting privacy definition, which can render the attacker powerless, is inherently strong.
Having formulated differential privacy from the hypothesis testing perspective, the next step is to specify the basis upon which the attacker determines which hypothesis to reject or accept. This necessitates a metric between the two probability distributions corresponding to the null and the alternative to quantify their difference. This motivates us to propose the following axiom as a self-evident requirement for any possible differential privacy definition.
@@
Formally, denote by $\mathrm{Proc}$ any post-processing operation on the output of the algorithm, such as dimension reduction or summarization. Axiom 2 states that
$D(M(S),M(S^{\prime}))\text{ and }D(\mathrm{Proc}(M(S)),\mathrm{Proc}(M(S^{\prime})))$
must be comparable and satisfy
@@ 2.3 A Representation Theorem
A proof of Theorem 3 is given in Section 2.4. For the construction of link functions for certain types of metrics $D$, see Proposition B.4 in [DRS22]. This theorem in the special case of $(\epsilon,\delta)$-differential privacy was first obtained by [KOV17].
@@ 1. Pure differential privacy
    is used to define pure differential privacy [DMNS06]. The corresponding link function is $d_{\infty}(f)=-\log(-f^{\prime}(1))$, where the derivative is understood as the left derivative. It is possible that $f^{\prime}(1)=0-$. Thus, $\epsilon=\infty$ is included for completeness.
@@ 2. Approximate differential privacy
    where $x_{+}$ denotes $\max\{x,0\}$. An algorithm $M$ satisfies $(\epsilon,\delta)$-differential privacy if
    $D_{\infty}^{\delta}(M(S),M(S^{\prime}))$
    is upper bounded by $\epsilon$ for any neighboring datasets $S$ and $S^{\prime}$. Smaller values of $\epsilon$ and $\delta$ provide stronger privacy protection. More precisely, $\delta$ should be set to a very small value, while a relatively large value of $\epsilon$ can still provide meaningful privacy guarantees [DKM+06, ND22].
@@
    where $f^{\ast}(x)=\sup_{y}(xy-f(y))$ is the convex conjugate of $f=T(P,Q)$. This representation is implied by a connection between $(\epsilon,\delta)$-differential privacy and hypothesis testing [WZ10, KOV17]. By varying $\delta$, all $(\epsilon,\delta)$-differential privacy guarantees taken together give the privacy profiles of the algorithm [BBG18], which are the dual representation of the trade-off function $f$.
@@ 2.4 The Blackwell Perspective
This theorem is a special case of a more general theorem called Blackwell’s informativeness theorem, which concerns the ranking of information structures and establishes an equivalence between rankings from different perspectives [Bla50]. The particular version presented in Theorem 4 is proven in [BG79] (Chapter 12, page 334). Its connection to differential privacy was first recognized by [KOV17] and further developed by [DRS22].
@@ 3.2 Achieving Differential Privacy
In [AV23], the authors called $Z$ a canonical noise distribution for $f$ if the three properties above are satisfied. Notably, a canonical noise distribution fully exploits the privacy budget conveyed by the trade-off function. This distribution admits an explicit construction based on the given trade-off function [AV23]. It has been extended to multivariate settings and shown to attain certain optimal statistical properties [AD22, AR23].
@@ 4.3 Shuffled Mechanism
The DAS utilizes a sequence of counting queries $\left\{\widebar{M}_{i}(S)\right\}_{i=1}^{m}$ with $m=72$, as specified in the privacy allocation file [US 22]. These queries cover 8 geographical levels, including nation, state, and county, with each level containing 9 queries. Each query is privatized as $M_{i}(S)=\widebar{M}_{i}(S)+\mathcal{N}_{\mathbb{Z}}(0,\sigma_{i}^{2})$. As each $\widebar{M}(S)$ is a counting query with sensitivity 1, the analysis of the privacy bound boils down to evaluating the trade-off function for testing the following hypotheses:
@@ 5 Concluding Remarks
In this review paper, we have explored differential privacy from a statistical viewpoint, motivated by the hypothesis testing framework. Drawing primarily from the work of [WZ10, KOV17, DRS22], we have shown that hypothesis testing is not merely a convenient tool but in fact the right language for reasoning about differential privacy. In light of this, $f$-differential privacy becomes a very natural definition that preserves both the hypothesis testing interpretation and the property of post-processing. Owing to its coherent mathematical structure, the $f$-differential privacy framework is accompanied by powerful tools, such as composition theorems, which enable tight and even lossless privacy analysis. This paper has demonstrated the effectiveness of this privacy definition in several important applications. Nevertheless, the review does not cover other notable applications, including federated learning, privacy auditing, and Markov chain Monte Carlo [ZCLS21, NHS+23, LMWR23].